OpenStudy (anonymous):

The velocity v of a soundwave varies jointly as the product of the frequency n and the wavelength ℓ. If the velocity of the wave is 12 ft/s when the frequency is 3 and the wavelength is 2.4, find the wavelength when the velocity of the soundwave is 18 ft/s and the frequency is 6. Round the solution to the nearest tenth.

OpenStudy (phi):

this is the same type of problem, velocity v varies jointly product of frequency n and the wavelength ℓ. that means v = k*n*l use the numbers to find k. after you find k, solve the 2nd part of the problem. can you find k ?

OpenStudy (anonymous):

3.6?

OpenStudy (phi):

for k ?

OpenStudy (phi):

can you show your work ?

OpenStudy (anonymous):

For k i got .83

OpenStudy (anonymous):

12 = k(3) (2.4) 12 = k(7.2) k = 1.67 I see what I did wrong

OpenStudy (anonymous):

18 = (1.67)(6)l l = 1.79

OpenStudy (phi):

yes, k = 5/3 exactly. so you should get 1.8 exactly

OpenStudy (anonymous):

Thank you phi

OpenStudy (phi):

yw